​(Individual

or component costs of

capital​)

Compute the cost of the​ following:

. Internal common equity when the current market price of the common stock is

​$45

The expected dividend this coming year should be

​$3.20

increasing thereafter at an annual growth rate of

8

percent. The​ corporation's tax rate is

33

percent.
Answer #1

Using Gordon Growth Model

Po = D1 / (Ke – g)

Where,

Po – Current share price = 45

D1 – Next year expected dividend = 3.20

Ke – Cost of equity = ?

G – Growth rate in dividend = 8%

45 = 3.20/ (Ke-.08)

Ke-.08 = 3.20/45

= 0.07111111111

Ke = 0.07111111111+.08

= 15.11%